OpenStudy (anonymous):

Use the double angle formula to show that: cos ( pi / 2) = 0 Thank you in advance!

OpenStudy (anonymous):

\[\cos\frac{\pi}{2}=\cos\left(2\cdot\frac{\pi}{4}\right)=\cos^2\frac{\pi}{4}-\sin^2\frac{\pi}{4}=\left(\frac{1}{\sqrt2}\right)^2-\left(\frac{1}{\sqrt2}\right)^2\]

OpenStudy (anonymous):

Thank you so much! I've been having so much difficulty with this.

OpenStudy (anonymous):

yw